OpenStudy (rcp031):

Andy takes 6 hours and Brinda takes 4 hours to read a book. Both started reading copies of the book at the same time. After how many hours is the number of pages to be read by Andy , twice that to be read by Brinda ? Assume Andy and Brinda read all the pages with constant pace. (A) 1 (B) 2 (C) 3 (D) 4

OpenStudy (rcp031):

Pls Help! @Zale101

OpenStudy (rcp031):

@Will.H

OpenStudy (rcp031):

@ShadowLegendX

OpenStudy (rcp031):

@agent0smith

OpenStudy (rcp031):

@KJSaif

OpenStudy (rcp031):

@zepdrix

OpenStudy (rcp031):

@inkyvoyd

OpenStudy (rcp031):

Am I correct in my approach, Let 'x' be the number of hours took andy to read twice that brinda Then number of pages read in an hour by Andy =1/6 of a book in an hour. Brinda= 1/4 of a book in an hour. Given question is . After how many hours is the number of pages to be read by Andy , twice that to be read by Brinda. (1/6)*x=2*(1/4) x=3 hours

OpenStudy (will.h):

No your answer is wrong if you subsitute x = 3 in the expression you made you will get \ 0.5 = 0.5 you said Andy should read twice Brinda that didn't happen here here they both read the same amount actually the whole question don't make any sense Let the number of pages x Andy takes 6 hrs so 6x Brinda takes 4 hrs so 4x 2(6x) = 4x 12x = 8x See if x = 1 will solve the equation (for the equation to be solved the left side must be twice the right side) 12 = 8 Since 12 is not twice the 8 then x = 1 is not the solution See if x = 2 will be the solution 12x = 8x 24 = 16 24 is not twice 16 so x= 2 is also not the solution See if x = 3 will be the solution 12x = 8x 36 = 24 also not the solution let’s see the last option 12x = 8x 48 = 32 None of the options makes sense And to me even the question doesn’t make sense.

OpenStudy (eliesaab):

none of the above should have been a choice and it would have been the right choice

OpenStudy (mrnood):

There is a phrase in the question that I htink is the source of the confusion: " to be read ". This means (IMO) thee number of pages LEFT to read After 3 hours Andy has read 3/6 and has 1/2 LEFT to read After 3 hours Brinda has read 3/4 and has 1/4 LEFT to read

OpenStudy (rcp031):

Thank you guys with your help finally i got the correct approach as well as answer.